# Ravenea: A Benchmark for Multimodal Retrieval-Augmented Visual Culture Understanding

### v1

@@ Abstract
As vision-language models (VLMs) become increasingly integrated into daily life, the need for accurate visual culture understanding is becoming critical. Yet, these models frequently fall short in interpreting cultural nuances effectively. Prior work has demonstrated the effectiveness of retrieval-augmented generation (RAG) in enhancing cultural understanding in text-only settings, while its application in multimodal scenarios remains underexplored. To bridge this gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a new benchmark designed to advance visual culture understanding through retrieval, focusing on two tasks: culture-focused visual question answering (cVQA) and culture-informed image captioning (cIC). Ravenea extends existing datasets by integrating over 10,000 Wikipedia documents curated and ranked by human annotators. With Ravenea, we train and evaluate seven multimodal retrievers for each image query, and measure the downstream impact of retrieval-augmented inputs across fourteen state-of-the-art VLMs. Our results show that lightweight VLMs, when augmented with culture-aware retrieval, outperform their non-augmented counterparts (by at least $3.2\%$ absolute on cVQA and $6.2\%$ absolute on cIC). This highlights the value of retrieval-augmented methods and culturally inclusive benchmarks for multimodal understanding.
@@ 1 Introduction
Vision-language models (VLMs) are increasingly deployed in real-world applications, from education to assistive technologies di2024educational ; de2024vqask ; bates-etal-2025-template ; yao2025countllm , where understanding not only visual content but also the surrounding cultural context is crucial. Despite achieving impressive performance on general tasks yue2024mmmu ; lin2014microsoft ; an2025multimodality ; an2025generalized , VLMs often struggle to capture cultural nuances, such as traditions, symbols, and region-specific practices that require external, culturally grounded knowledge (nayak-etal-2024-benchmarking, ; liu2025culturevlm, ; khanuja-etal-2024-image, ; cao2024exploring, ). For example, as shown in Figure 1, a VLM may incorrectly identify the season of a festival scene as ‘Autumn’, overlooking that the image depicts Kyoto’s Gion Festival, which occurs in July and corresponds to ‘Summer’. A promising approach to address this limitation is the integration of external knowledge through retrieval-augmented generation (RAG) lewis2020retrieval , which has shown success in improving cultural awareness in language models (seo2025valuesrag, ; lertvittayakumjorn2025towards, ). However, prior work in the culture domain has predominantly been confined to text-only settings. Meanwhile, existing culture-related multimodal datasets primarily focus on evaluating VLM outputs on culturally oriented tasks, with limited emphasis on the integration of external cultural knowledge. As a result, the potential of RAG to improve multimodal cultural understanding remains underexplored.
Figure 1: Effectiveness of culture-aware RAG. Given a culturally grounded visual question, VLMs enhanced with culture-aware RAG—retrieving relevant Wikipedia documents—generate more accurate answers than their non-RAG counterparts. (see Section 5.2).
@@
To bridge the gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a manually curated dataset designed to evaluate cultural understanding in VLMs with retrieval support. We construct Ravenea based on two existing datasets: CVQA (romero2025cvqa, ), which includes culturally relevant visual questions and corresponding answers, and CCUB (liu2023towards, ), offering culturally contextualized captions to foster inclusivity in text-to-image generation  [^1] We reuse the cultural captions from CCUB as ground-truth references for the inverse task, image-to-text generation, specifically for culture-aware image captioning.. For each instance drawn from the source datasets, we append a set of Wikipedia documents that have been human-ranked based on their cultural relevance to the associated image. This curation effort, designed to ensure broad cultural representation, contains data related to eight countries and spans eleven diverse categories, comprising more than 1,800 images and 10,000 human-ranked documents. Ravenea thus provides a retrieval-augmented benchmark for evaluating cultural sensitivity in multimodal retrievers, and further allows for assessing how well VLMs integrate and make use of retrieved cultural context. Specifically, we focus on two culturally grounded tasks: (i) culture-focused visual question answering (cVQA) and (ii) culture-informed image captioning (cIC). We select these two tasks for their centrality in assessing cultural understanding in VLMs—question answering tests context-aware reasoning, while captioning evaluates generation sensitivity to cultural nuances.
@@
With Ravenea, we first train and evaluate seven multimodal retrievers that use both visual and textual inputs to retrieve Wikipedia documents for a given query image based on their cultural relevance. Then, we evaluate a diverse set of state-of-the-art (SOTA) VLMs, including GPT-4.1 achiam2023gpt , LLaVA-OneVision-7B li2025llavaonevision , Pixtral agrawal2024pixtral , Phi-4 Multimodal abouelenin2025phi , the Gemma3 family team2025gemma , the Qwen2.5-VL family bai2025qwen2 , the InternVL3 family zhu2025internvl3 , and the Deepseek-VL2 family wu2024deepseek , each with and without multimodal retrieval, to assess the impact of retrieval augmentation on cultural understanding. Our dataset provides a testbed for assessing the cultural relevance capabilities of multimodal retrievers and the effectiveness of VLMs in consuming and using such retrieved cultural context.
@@
  • Ravenea benchmark: We introduce Ravenea, the first benchmark aimed at evaluating VLMs and multimodal retrieval in leveraging external knowledge for visual culture understanding. It comprises a diverse, large-scale collection of human-curated, culturally related documents linked to images from eight countries across eleven categories, enabling evaluation on two tasks: culture-focused visual question answering and culture-informed image captioning. (Section 3)
@@
  • Cultural grounding annotations enhance multi-modal retrieval: We evaluate seven retrievers that integrate visual and textual cues to retrieve culturally relevant documents. We find fine-tuning retrievers on culture-targeted annotations leads to marked gains in retrieval accuracy, highlighting the value of explicit cultural supervision. (Section 6.3)
@@
  • Benefits of culture-aware retrieval: Culture-aware retrieval boosts task performance across VLMs, with lightweight models showing the greatest improvement. This suggests that such retrieval can seamlessly integrate into downstream VLM tasks, enhancing their performance. (Section 5.2)
@@
  • Cross-cultural variation: Evaluation across eight countries reveals that VLMs exhibit distinct cultural preferences, with each model favoring different regional contexts—suggesting model-specific cultural biases. (Section 6.2)
@@
Figure 2: Ravenea: A Multimodal Retrieval-Augmented Visual culturE uNdErstAnding dataset. Left: Examples of cVQA and cIC tasks. Middle: Geographic and categorical distribution of cultural references. Right: Performance comparison of 14 VLMs, evaluated with and without integration of our culture-aware retriever. Here, CaCLIP="culture-aware CLIP-L/14@224px ".
2 Related Work
Retrieval augmentation for cultural understanding. Retrieval augmentation has demonstrated significant efficacy for culture-related NLP tasks (conia-etal-2024-towards, ; hu-etal-2024-bridging, ; fung-etal-2023-normsage, ). Prior work uses sources like the World Values Survey for cultural question answering seo2025valuesrag ; yadav-etal-2025-beyond , or retrieves web and knowledge base content to enhance cultural contextualization lertvittayakumjorn2025towards . In multimodal settings, however, cultural retrieval remains underexplored. While some approaches retrieve culture-relevant images to fine-tune VLMs liu2025culturevlm , they require additional training. In contrast, our method introduces a plug-and-play retrieval system that enriches cultural grounding at inference time without modifying the base model.
Vision-language culture datasets. Several recent studies examine cultural understanding in VLMs through multicultural VQA liu-etal-2021-visually ; nayak-etal-2024-benchmarking , cuisine recognition winata-etal-2025-worldcuisines ; li-etal-2024-foodieqa , and concept-based image retrieval bhatia-etal-2024-local . Others benchmark cultural entity recognition using Wikipedia-based prompts nikandrou-etal-2025-crope or curate culture-specific image sets for value-based tasks yadav-etal-2025-beyond . While these efforts highlight current VLM limitations, they often rely on static, manually curated datasets. A natural extension to address these deficiencies is through culture-relevant multimodal retrieval—an area lacking dedicated datasets. Our work specifically aims to fill this critical research gap.
3 Ravenea Dataset
We construct our dataset by building culturally relevant document lists for each image in two existing culture-grounded datasets: CVQA romero2025cvqa , a widely used dataset for culture-focused VQA, and CCUB liu2023towards , a dataset designed to mitigate cultural bias in text-to-image generation with culturally contextualized captions. To ensure broad geographic and cultural coverage, we curate a diverse subset comprising images from seven countries in CVQA: China, Nigeria, Russia, Spain, Mexico, India, and Indonesia, and all five countries in CCUB: China, Korea, India, Mexico, and Nigeria (see Figure 2). After that, we separate the dataset construction process into three critical stages shown in Figure 3.
3.1 Dataset Construction
Data collection. The data collection process consists of two main steps: (i) culture-related captioning and (ii) document retrieval. For culture-related captioning, we generate culturally grounded captions for each image to facilitate more effective attachment of relevant documents. Since the CVQA lacks captions and the CCUB provides only brief descriptions, we employ GPT-4o to generate richer, culturally informative captions (see the prompt example in Table 10). For document retrieval, we first conduct a coarse filtering using the generated cultural captions as queries for a BM25 Robertson2009ThePR retriever to extract the semantically relevant documents from a large-scale corpus comprising over six million English Wikipedia documents  [^2] https://huggingface.co/datasets/wikimedia/wikipedia. To mitigate the impact of inaccurate captions and ensure precise document relevance, we then perform human annotation on the retrieved documents.
@@ 3.1 Dataset Construction
Relevance annotation. Based on the initial BM25 retrieval results, we refine the cultural relevance label of retrieved documents via human annotation. For each image-caption pair, annotators are presented with the top 10 Wikipedia documents retrieved by BM25. They are asked to assess whether each document provides meaningful background or contextual information that is relevant to the culture described in the caption or the image (see Appendix K). Specifically, we decomposed cultural relevance into three interpretable and independently verifiable dimensions: Country association: Is the topic of the Wikipedia article associated with the same country as the image and its caption? Topic alignment: Does the topic of the Wikipedia article align with the semantic category of the image and its caption? Explicit visual representation: Is the topic of the Wikipedia article explicitly mentioned or visually represented in the image and its caption? Each dimension is framed as a binary (True / False) question to reduce ambiguity and improve annotation consistency. However, for the country association dimension (the first listed), we introduce an additional label, "Cannot be determined", to handle cases where this association is unclear from the annotator’s perspective. Additionally, annotators are also instructed to include the title and URL of any relevant Wikipedia article they believe is missing from the top-10 retrieved results. These manually suggested articles are treated as the most cultural references closely related to the given image (see details in Appendix G).
Quality control. To ensure the quality and consistency of our annotations, we implement several quality control methods. First of all, prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines.
Table 1: Statistics of the Ravenea dataset. The dataset is constructed by curating existing sources and augmenting them with over 10,000 wiki-derived documents to broaden cultural knowledge coverage and enhance content diversity.
Dataset Images Documents Pairs Questions Captions
CVQA 1,213 8,319 12,130 2,331 -
CCUB 655 4,441 6,550 - 655
Ravenea 1,868 11,580 18,680 2,331 655
To ensure proper understanding of the guidelines, annotators are required to complete a mock annotation test and correctly answer all questions before proceeding with the actual annotation tasks (see Figure 14). We also perform an additional quality check on a subset of the dataset. Specifically, for each selected countries, we employ an additional local quality checker who is tasked with manually reviewing the annotations to verify their accuracy and adherence to the guidelines. The quality checker reviews a random sample of annotated items, focusing on both the relevance labels and the justification behind any edge cases, such as borderline relevance or use of the “Cannot be determined” label. If inconsistencies or deviations from the annotation guidelines are identified, the affected samples are flagged for re-annotation. The overall acceptance rate from the meta quality checkers is 98.2%. The inter-annotator agreement (IAA) Cohen’s Kappa ($\kappa$artstein2017inter between the meta checker and annotator on the sampled annotations is 0.83.
Figure 4: Examples demonstrating the impact of CaCLIP Wikipedia retrieval integration on cVQA and cIC tasks using DeepseekVL2-Tiny. When augmented with culture-aware retrieval, the model exhibits enhanced sensitivity to cultural context.
3.2 Dataset Statistics
We present the statistics of our dataset in Table 1. The dataset comprises a total of 1,868 culturally diverse images, with approximately 65% originating from CVQA and the remaining 35% from CCUB. Each image is paired with a GPT-4o-generated cultural caption, as well as the top-10 ranked Wikipedia documents retrieved via a cultural relevance scoring pipeline (Figure 3), yielding a total of 18,680 image-document pairs (see illustrated examples in Figure 4). The collection spans eight countries and encompasses a broad spectrum of cultural domains, such as traditional attire, festivals, architecture, cuisine, and social practices (see more statistics in Appendix E).
4 Culture-aware Multimodal Retriever
Leveraging the Ravenea dataset, we train and evaluate seven multimodal retrievers to retrieve culturally relevant Wikipedia documents using both visual and textual inputs. We fine-tune five representative models—spanning both generative and discriminative paradigms—to optimize multimodal document retrieval. Performance is evaluated using standard retrieval metrics, including Mean Reciprocal Rank (MRR) shi2012climf , Precision@k (P@k) jarvelin2017ir , and Normalized Discounted Cumulative Gain (nDCG@k) wang2013theoretical , where $k\in\{1,3,5\}$. We integrate responses from three annotation questions per data point into a continuous scale ranging from $-3$ to $3$, where higher values indicate stronger cultural relevance. We fine-tune a VisualBERT-based li2019visualbert ; li-etal-2020-bert-vision reranker following standard BERT-style setups nogueira2019passage , and adapt two multimodal generators—VL-T5 cho2021unifying and LLaVA-OneVision-7B li2025llavaonevision —for end-to-end document retrieval chen2022corpusbrain ; yuan2024asking ; sun2023learning . To enhance cultural awareness in the contrastive retrieval, we introduce Culture-Aware Contrastive (CAC) learning, a supervised learning framework compatible with both CLIP and SigLIP architectures. We denote the culture-aware fine-tuned versions of CLIP-L/14@224px and SigLIP2-SO/14@384px using CAC as CaCLIP and CaSigLIP2, respectively.

### v2

@@ Abstract
As vision-language models (VLMs) become increasingly integrated into daily life, the need for accurate visual culture understanding is becoming critical. Yet, these models frequently fall short in interpreting cultural nuances effectively. Prior work has demonstrated the effectiveness of retrieval-augmented generation (RAG) in enhancing cultural understanding in text-only settings, while its application in multimodal scenarios remains underexplored. To bridge this gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a new benchmark designed to advance visual culture understanding through retrieval, focusing on two tasks: culture-centric visual question answering (cVQA) and culture-informed image captioning (cIC). Ravenea extends existing datasets by integrating 11,396 unique Wikipedia documents curated and ranked by human annotators. Through the extensive evaluation on seven multimodal retrievers and seventeen VLMs, Ravenea reveals some undiscovered findings: (i) In general, cultural grounding annotations can enhance multimodal retrieval and corresponding downstream tasks. (ii) VLMs, when augmented with culture-aware retrieval, generally outperform their non-augmented counterparts (by averaging $6\%$ on cVQA and $11\%$ on cIC). (iii) Performance of culture-aware retrieval augmented varies widely across countries. These findings highlight the limitations of current multimodal retrievers and VLMs, underscoring the need to enhance visual culture understanding within RAG systems. We believe Ravenea offers a valuable resource for advancing research on retrieval-augmented visual culture understanding.
Project page: jiaangli.github.io/ravenea
@@ 1 Introduction
Vision-language models (VLMs) are increasingly deployed in real-world applications, from education to assistive technologies (Baker et al., 2021; Di Nuovo et al., 2024; De Marsico et al., 2024; Karamolegkou et al., 2025), where understanding not only visual content but also the surrounding cultural context is crucial. Despite achieving impressive performance on general tasks (Lin et al., 2014; Yue et al., 2024; Bai et al., 2025a), VLMs often struggle to capture cultural nuances, such as traditions, symbols, and region-specific practices that require external, culturally grounded knowledge (Nayak et al., 2024; Khanuja et al., 2024; Liu et al., 2025; Winata et al., 2025). For example, when analyzing images of cultural events, VLMs may misinterpret region-specific attire as generic clothing, thereby overlooking its ceremonial significance (Nayak et al., 2024). What’s more, they also risk reinforcing existing cultural biases, as biased collections or metadata frequently lead VLMs to perform much better in mainstream culture contexts (e.g. western culture) while marginalizing minority traditions (Nayak et al., 2024). A promising approach to mitigate the limitations is the integration of external knowledge through retrieval-augmented generation (RAG) (Lewis et al., 2020), which has shown success in improving cultural awareness in language models (Seo et al., 2025; Lertvittayakumjorn et al., 2025). However, prior work on culture has largely remained in text-only settings, while existing multimodal culture-related datasets mostly probe VLMs’ memorized cultural knowledge instead of testing their understanding of culture in real-world contexts. Consequently, it remains unclear to what extent current mutlimodal retrievers support reliable cultural retrieval, and whether RAG can effectively advance multimodal cultural understanding for VLMs. This shortfall underscores the need for a dedicated benchmark that probes various facets of visual culture comprehension for multimodal retrievers and VLMs.
Figure 1: Samples from Ravenea. Ravenea evaluates the multimodal culture-awareness RAG with VLMs across eight countries spanning four continents: Asia, Africa, Europe, and Latin America. It comprises two tasks designed to assess visual culture understanding abilities and the effect of extra cultural knowledge across different countries. With 1,868 instances and 11,369 human-ranked documents, Ravenea covers eleven real-world scenarios, including architecture, cuisine, art, etc.
@@
To bridge the gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a manually curated dataset designed to evaluate multicultural understanding in VLMs with retrieval support. We construct Ravenea based on two existing datasets: CVQA (Romero et al., 2025), which includes culturally relevant visual questions and corresponding answers, and CCUB (Liu et al., 2023), offering culturally contextualized captions to foster inclusivity in text-to-image generation. [^1] We reuse the cultural captions from CCUB as ground-truth references for the inverse task, image-to-text generation, specifically for culture-aware image captioning. For each instance drawn from the source datasets, we append a set of Wikipedia documents that have been human-ranked based on their cultural relevance to the associated image. This curation effort, designed to ensure broad cultural representation, contains data related to eight countries and spans eleven diverse categories (e.g. architecture, art, cuisine, etc), comprising 18,680 human-labeled image-document pairs. Ravenea thus provides a multimodal retrieval-augmented benchmark for evaluating multicultural sensitivity of various retrievers, and further allows for assessing how well VLMs utilize retrieved cultural context in their reasoning process. Specifically, as shown in Figure 1, we focus on two culturally grounded tasks: (i) culture-centric visual question answering (cVQA) and (ii) culture-informed image captioning (cIC).
@@
With Ravenea, we first train and evaluate seven multimodal retrievers that use both visual and textual inputs to retrieve Wikipedia documents for a given image query based on their cultural relevance. We then evaluate 17 widely used VLMs, both with and without multimodal retrieval, spanning open-source and proprietary models as well as sizes from 2B to 78B parameters, to assess the effect of retrieval augmentation on cultural understanding in the cVQA and cIC tasks. Our benchmark provides a testbed for evaluating the cultural relevance capabilities of multimodal retrievers and the effectiveness of VLMs in consuming and using such retrieved cultural context.
@@
  • Ravenea benchmark: We introduce Ravenea, the first benchmark for evaluating VLMs in using external knowledge for visual culture understanding. It covers eight countries across eleven categories, linking images to human-ranked Wikipedia documents on two tasks: cVQA and cIC.
@@
  • Cultural grounding annotations enhance multi-modal retrieval: We evaluate seven retrievers that integrate visual and textual cues to retrieve culturally relevant documents. We find fine-tuning retrievers on culture-targeted annotations leads to marked gains in retrieval accuracy, highlighting the value of explicit cultural supervision.
@@
  • Benefits of culture-aware retrieval: Culture-aware retrieval boosts task performance across VLMs, with lightweight models showing the better improvement. This suggests that such retrieval can seamlessly integrate into downstream VLM tasks, enhancing their performance.
@@
  • Cross-cultural variation: Evaluation across eight countries reveals that VLMs, both with and without cultural RAG exhibit distinct cultural preferences, with each model favoring different regional contexts, suggesting model-specific cultural biases.
2 Ravenea Construction
@@
Figure 2: Ravenea: A Multimodal Retrieval-Augmented Visual culturE uNdErstAnding dataset. Geographic and categorical distribution of cultural references in Ravenea.
We present Ravenea, the first benchmark explicitly designed to assess how retrieval augmentation influences the visual cultural understanding capabilities of VLMs. It is derived originally from 30 countries; see Table 4. Each image is paired with 10 algorithm-ranked documents based on cultural relevance. To ensure high data quality while maximizing linguistic diversity across major language families, we sample data from eight countries spanning multiple continents: China, Nigeria, Russia, Spain, Mexico, India, Indonesia, and Korea. [^2] List of languages by total number of speakers. We then apply human re-ranking–based annotation to the collected documents to construct Ravenea. Consequently, Ravenea comprises 1,868 instances, including 2,331 image-text questions, 655 image-caption pairs, and 18,680 image-document pairs, supporting two tasks: cVQA and cIC, respectively.
2.1 Taxonomy
Motivated by the increasing need for VLMs to understand cultural context (Yadav et al., 2025a), we categorize the cultural understanding capability into two core tasks and several detailed axes, as shown in Figure 2. Additional dataset statistics are provided in Appendix D.
cVQA.   This task evaluates a model’s ability to perceive and reason over culturally grounded visual content, often requiring external, culturally specific knowledge beyond the image itself. Building on the CVQA dataset (Romero et al., 2025), which contains culture-related QA pairs, we augment it with culturally relevant documents that supply the additional knowledge required for RAG evaluation. Therefore, each instance contains an image, a question, top 10 human-ranked documents based on culture relevance, and multiple-choice options, with only one correct answer.
cIC.   This task assesses the VLM’s ability to generate captions that are sensitive to and incorporate cultural nuances. It evaluates the model’s generation sensitivity by requiring it to produce culturally contextualized captions for a given image, leveraging information from human-curated, culturally related documents. The task reuses cultural captions from the CCUB dataset (Liu et al., 2023) as ground-truth references for this task. As a result, for cIC, each instance consists of an image, a human-written caption, and top 10 culturally-related documents.
@@ 2.1 Taxonomy
2.2 Dataset Curation
Data collection.   As shown in Figure 3, the data collection process consists of two main steps: (i) culture-related captioning and (ii) document retrieval. For culture-related captioning, we generate culturally grounded captions for each image to facilitate more effective attachment of relevant documents. Since the CVQA lacks captions and the CCUB provides only brief descriptions, we employ GPT-4o [^3] Knowledge cutoff: Oct 01, 2023. (Achiam et al., 2023) to generate richer, culturally informative captions (see the prompt example in Table 13). [^4] We used LLMs to generate culture-informed captions solely as auxiliary references for annotators; these LLM outputs were not directly included in the final dataset. For document retrieval, we begin with a coarse-grained filtering stage in which the generated cultural captions serve as queries to a BM25 retriever (Robertson and Zaragoza, 2009). This stage retrieves a pool of semantically relevant candidates from a large-scale corpus of over six million English Wikipedia articles. [^5] Wikimedia/wikipedia To mitigate the impact of inaccurate captions and ensure precise document relevance, we then perform human annotation on the retrieved documents.
Cultural relevance annotation.   Based on the initial BM25 retrieval results, we refine the cultural relevance label of retrieved documents via human annotation. For each image-caption pair, annotators are presented with the top 10 Wikipedia documents retrieved by BM25. They are asked to assess whether each document provides background or information that is relevant to the culture described in the caption or the image (see interfaces in Appendix L). Specifically, we decomposed cultural relevance into three interpretable and independently verifiable dimensions:
Country association: Is the topic of the Wikipedia article associated with the same country as the image and its caption? Topic alignment: Does the topic of the Wikipedia article align with the semantic category of the image and its caption? Explicit visual representation: Is the topic of the Wikipedia article explicitly mentioned or visually represented in the image and its caption?
@@ 2.2 Dataset Curation
Each dimension is framed as a binary (True / False) question to reduce ambiguity and improve annotation consistency. However, for the country association dimension (the first listed), we introduce an additional label, "Cannot be determined", to handle cases where this association is unclear from the annotator’s perspective. Additionally, annotators are also instructed to include the title and URL of any relevant Wikipedia article they believe is missing from the top-10 retrieved results. These manually suggested articles are treated as the most culturally relevant references for the given image. Detailed annotation statistics, alongside positive, negative and missing ratios by countries, are presented in Appendix E.
Quality control.   For caption quality, we employ two independent local annotators assessed both objective accuracy and subjective quality (naturalness and completeness) in a subset, achieving 92% and 94% accuracy with an inter-annotator agreement (IAA) Cohen’s Kappa ($\kappa$(Artstein, 2017) of 0.85, and yielding high subjective ratings (naturalness: 4.98/5, completeness: 4.66/5). The results demonstrate that the captions are reliable for downstream retrieval of Wikipedia documents. For annotation quality, a separate group of annotators were trained with detailed guidelines and a mock test before performing the actual annotation. A meta quality checker further validated a random subset, achieving a 98.2% acceptance rate. The IAA Cohen’s Kappa ($\kappa$) between the meta checker and annotator on the sampled annotations is 0.83. More details are presented in Appendix F.
3 Experimental Setup
With Ravenea, we train and evaluate seven multimodal retrievers to retrieve culturally relevant Wikipedia documents using multimodal inputs, spanning both generative and discriminative paradigms. To enhance cultural awareness in the contrastive retrieval, we introduce Culture-Aware Contrastive (CAC) learning (details in Section 3.2), a supervised learning framework compatible with both CLIP (Radford et al., 2021) and SigLIP (Tschannen et al., 2025) architectures. Then, we evaluate the effectiveness of these retrievers with 17 widely used VLMs across the cVQA and cIC downstream tasks. More details are in Appendix I.
3.1 Evaluation Metrics
For multimodal retrieval, the performance is evaluated using standard retrieval metrics, including Mean Reciprocal Rank (MRR) (Shi et al., 2012), Precision@k (P@k) (Järvelin and Kekäläinen, 2017), and Normalized Discounted Cumulative Gain (nDCG@k) (Wang et al., 2013), where $k\in\{1,3,5\}$.
Table 1: Kendall’s $\tau$ rank correlation (Kendall, 1938) between automatic metrics and human judgments for the cIC task. Statistically significant correlations ($p<0.05$) are marked with ✓. Our proposed metrics correlate stronger with human evaluation than the others.
Rouge-L CIDER BERTScore CLIPScore RegionScore (ours)
-0.172 ✗ -0.316 ✗ -0.011 ✗ 0.139 ✗ 0.442 ✓
For cVQA, we use accuracy as the evaluation metric, as all tasks follow a multiple-choice format. Accuracy is computed per VLM and per retriever, representing the proportion of correctly answered questions. For cIC, we employ several metrics for evaluating the generated caption, including ROUGE-L (Lin, 2004), CIDEr (Vedantam et al., 2015), BERTScore (Zhang* et al., 2020), and CLIPScore (Hessel et al., 2021).
To further evaluate the cultural relevance of the generated captions, we conducted a human study in which four researchers selected the most accurate captions from 14 VLMs. [^6] The Qwen3-VL models were omitted from the human evaluation for the cIC task because the used model variants were not released at that time. We find a significant mismatch between automatic metrics and human judgments of cultural appropriateness (Table 1). To bridge this gap, we further introduce RegionScore, a novel evaluation metric designed to quantify the extent to which captions reference specific geopolitical regions. See details in Appendix G and H.
RegionScore.   Let ${\mathbb{I}}=\{I_{1},I_{2},\dots,I_{n}\}$ denote the set of images. For each image $I_{i}$, let ${\bm{g}}^{(i)}$ denote its generated caption, and let ${\bm{c}}^{(i)}$ represent the country name associated with $I_{i}$. We define ${\mathbb{A}}_{i}$ as the set of adjectives or demonyms corresponding to ${\bm{c}}^{(i)}$. For convenience, we introduce the set of region-related terms associated with $I_{i}$: ${\mathbb{T}}_{I_{i}}=\{{\bm{c}}^{(i)}\}\cup{\mathbb{A}}_{i}$. The instance-level $R(\displaystyle{\bm{g}}^{(i)},I_{i})$ metric for sample $I_{i}$ and ${\bm{g}}^{(i)}$ is defined as:
[…]
The overall RegionScore for predicted captions are then computed as:
$RegionScore_{\displaystyle{\mathbb{I}}}=\frac{1}{|\displaystyle{\mathbb{I}}|}\sum_{I_{i}\in\displaystyle{\mathbb{I}}}R(\displaystyle{\bm{g}}^{(i)},I_{i}),\quad RegionScore_{\displaystyle{\mathbb{I}}}\in[0,1]$ (2)
RegionScore is designed to complement, rather than replace, existing general-purpose captioning metrics (e.g., CLIPScore), which primarily assess semantic alignment but do not explicitly capture geographic specificity. Therefore, RegionScore serves as a proxy for regional awareness by quantifying the proportion of captions that include explicit references of a country or its common demonyms, with higher values indicating stronger region awareness. In CCUB, $\text{RegionScore}_{\text{ground truth}}$ is 99%.